OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU A MEDAL AND FAN What is the product in lowest terms? 6 3 ~ -- ∗ -- 11 4 A) 9 ~ --- 22 B) 6 ~ --- 13 C) 9 --- 22 D) 6 --- 13 (IMAGINE THE ~ SIGN AS A NEGATIVE SIGN)

OpenStudy (ccswims):

Well first you multiply the numerators \[6 \times 3\]\[= 18\] Then you multiply the denominators \[-11\times4\]\[= -44\]

OpenStudy (ccswims):

So now we have \[\frac{ -18 }{ 44 }\]

OpenStudy (ccswims):

But we can simplify that by 2

OpenStudy (ccswims):

\[\frac{ -9 }{ 22 }\]

OpenStudy (ccswims):

And it must be a negative because whenever you multiply/divide a negative with a positive the outcome is always negative.

OpenStudy (winsteria):

So it's -9/22?

OpenStudy (ccswims):

ya

OpenStudy (winsteria):

thank you

OpenStudy (anonymous):

hey there best friend

OpenStudy (anonymous):

ok.... i guess your not in happy mood?

OpenStudy (winsteria):

what?

OpenStudy (anonymous):

xD i was saying hi